# Reinforcement Learning with Intrinsically Motivated Feedback Graph for Lost-sales Inventory Control

### v1

@@ 5 Conclusion
dresses the challenge of sample inefficiency for RL methods in lost-sales IC problems. We design a novel decision framework integrating RLFG and IME to boost the sample efficiency of RL methods. We first specially tailored FG only based on the general property of lost-sales IC problems rather than strong heuristics to generate side experiences to aid RL update. We then conduct a theoretical analysis to demonstrate our method’s effectiveness with Q-learning as an example. The analysis shows that FG decreases the sample complexity by improving the update probabilities for all state-action pairs. Additionally, we design an intrinsic reward to fully utilize FG for lost-sales IC problems based on the analysis result. Experimental results show that our approach greatly enhances RL’s sampling efficiency in IC problems, which is consistent with the theoretical analysis. As for the limitation, we only analyze the magnitude relationship between $\mu_{\text{min}}$ and $\tilde{\mu}_{\text{min}}$; a detailed multiplicative relationship still needs to be proved.

### v2

@@ 5 Conclusion
This paper addresses the challenge of sample scarcity for RL methods in lost-sales IC problems. We design a novel training framework integrating RLFG and IME to boost the sample efficiency of RL methods. We first specially tailored FG using general lost-sales properties, environment-specific heuristics, to generate side experiences to aid RL updates. We then conduct a theoretical analysis to demonstrate our method’s effectiveness with Q-learning as an example. The analysis shows that FG decreases the sample complexity by improving the update probabilities for all state-action pairs. Additionally, we design an intrinsic reward to fully utilize FG for lost-sales IC problems based on the theoretical insights. To validate the performance and generalization ability of our proposed method, we have done experiments on single-item, multi-item, and multi-echelon lost-sales IC environments. Experimental results show that our approach greatly enhances RL’s sampling efficiency in IC problems, which is consistent with the theoretical analysis.